OpenStudy (anonymous):

calculate double integral xy^2/x^2+1 where R ={(x,y)/0<=x<=1,-3<=y<=3}

sam (.sam.):

\[\int\limits_{}^{}\int\limits \frac{x y^2}{x^2+1} \, dxdy\]??????

OpenStudy (anonymous):

limits of x=[0,1] and y=[-3,3]

sam (.sam.):

do the inner first \[\int\limits \frac{x y^2}{x^2+1} \, dx\]

sam (.sam.):

\[\begin{array}{l} \text{Factor out constants:} \\ \text{}=y^2\int\limits \frac{x}{x^2+1} \, dx \\\end{array}\]

sam (.sam.):

u-sub u=x^2+1 du=2xdx

sam (.sam.):

\[\frac{y^2}{2}\int\limits \frac{1}{u} \, du\] \[ [ \frac{1}{2} y^2 \ln (u)]_{}^{}\] \[\frac{1}{2} y^2 \ln \left(x^2+1\right)|_{0}^{1}\]

sam (.sam.):

\[\int\limits_{-3}^{3}\frac{1}{2} y^2 \ln \left(2\right)dy\]

sam (.sam.):

\[\int\limits \frac{1}{2} y^2 \log (2) \, dy\] \[\frac{1}{6} y^3 \log (2)|_{-3}^{3}\]

sam (.sam.):

Solve

OpenStudy (anonymous):

i am getting \[\int\limits_{-3}^{3} y^2/2 (\ln(2) -\ln (1))dy\]

sam (.sam.):

There's some mistake in your calculations? or maybe its inverted \[\int\limits\limits_{}^{}\int\limits\limits \frac{x y^2}{x^2+1} \, dxdy \leftarrow \rightarrow \int\limits\limits_{}^{}\int\limits\limits \frac{x y^2}{x^2+1} \, dydx\]

OpenStudy (anonymous):

final anwer is 9 ln2/ln1 ?

OpenStudy (anonymous):

i meant 9ln2

sam (.sam.):

yes

OpenStudy (anonymous):

thanx